# vCache: Verified Semantic Prompt Caching

### v4

@@ Appendix A Notation Glossary
Appendix D Evaluation Results
D.1 SemCacheSearchQueries Benchmark
@@
Figure 7: vCache meets the user-defined maximum error rate bound $\delta$ with steadily increasing cache hit rates (vCache is learning). GPTCache shows increasing error and hit rates, illustrating the unreliability of static thresholds. The static baseline uses a fixed threshold of 0.985. See Figure 8 for a threshold vs. $\delta$ Pareto comparison.
@@
We evaluate whether vCache satisfies the user-defined error rate $\delta$ while maintaining competitive performance. As shown in Figure 7 (left), vCache consistently remains below the specified error bound across all tested $\delta$ values. Moreover, as the error rate stabilizes, vCache continues to improve its cache hit rate (Figure 4, right). In contrast, GPTCache exhibits increasing error rates as the sample size grows, despite improved hit rates. This trend reflects a fundamental limitation of static thresholds: maintaining a bounded error rate requires continuously increasing the threshold. Over time, no static threshold below 1.0 may suffice to satisfy a strict error constraint, making such systems difficult to tune and unreliable at scale.
Figure 8: Pareto comparison across a range of thresholds and $\delta$ values on 150k samples from the SemCacheSearchQueries benchmark. While vCache outperforms static baselines while its error rate remains bounded. In contrast, static-threshold methods like GPTCache require increasing thresholds to maintain low error (i.e., a threshold of 0.99 for a 1.7% error rate after 150k samples), shifting the Pareto curve upward over time. This highlights the limitations of static thresholds in sustaining error-rate guarantees as sample size grows.
@@
D.2 Embedding Generation Overhead
@@
Figure 9: Embedding computation latency distributions across models, shown as 95th percentile whisker plots. GTE_FT = GteLargeENv1-5 (Zhang et al., 2024) fine-tuned (Zhu et al., 2024). E5 Large v2 FT = E5-large-v2 (Wang et al., 2022) fine-tuned (Zhu et al., 2024)
@@
D.3 Logistic Regression Latency Overhead
@@
D.4 vCache in combination with vLLM
@@
Baseline Config
Overall
Duration
Avg. LLM
Inference Latency
Avg. Emb.
Latency
Cache
Hit Rate
Error
Rate
@@
vLLM
+ vCache
$\delta=0.01$ 214 min 0.32 sec 0.018 sec 18.1% 0.4%
@@
D.5 Additional Baseline Evaluations
@@
Baseline
Threshold/
Delta
Error
Rate
Cache hit
Rate
Comments and Observations
@@ D.6 Threshold Dilemma
Appendix E Benchmark Creation
To the best of our knowledge, no open-source benchmarks currently exist for evaluating the performance and applicability of semantic caching systems. To address this gap, we construct and release three diverse benchmarks [^3] https://huggingface.co/vCache, each designed to reflect a distinct real-world use case: classification tasks, conversational chatbots, and search engines. This section describes the motivation and construction process behind each benchmark.
E.1 SemCacheClassification Benchmark
@@ E.3 SemCacheSearchQueries Benchmark
Appendix F Hyperparameter Guidance

### v5

@@ Appendix A Notation Glossary
Appendix D Convergence Speed Estimation
For each cached prompt $y=\mathrm{nn}(x)$, vCache fits the sigmoid model $\mathcal{L}(s(x),t,\gamma)$ to the meta-data $\mathcal{O}(y)$. Under the standard regularity assumptions for logistic regression, the MLE $(\hat{t}_{y},\hat{\gamma}_{y})$ converges to the true embedding-specific parameters $(t_{y},\gamma_{y})$ at the usual parametric rate $\mathcal{O}(1/\sqrt{n_{y}})$, where $n_{y}$ denotes the number of explored queries whose nearest neighbor is $y$. Any embedding-specific threshold derived from this model (e.g., the similarity at which $\Pr(c(x)=1\mid x,\mathcal{D})$ exceeds $1-\delta$, or the required exploration probability $\tau_{\mathrm{nn}(x)}(s(x))$) is a smooth function of $(t_{y},\gamma_{y})$ and therefore inherits the same $\mathcal{O}(1/\sqrt{n_{y}})$ convergence rate via the delta method. Intuitively, as vCache collects more labels for a given neighbor, the decision boundary $t$ and the corresponding policy $P_{\text{vCache}}(s(x),\mathcal{O}(\mathrm{nn}(x)),\delta)$ stabilize at this standard parametric speed.
Appendix E Empirical Justification of the Sigmoid Model
We model, for each cached embedding $nn(x)$, the probability $\Pr(c(x)=1\mid x,\mathcal{D})$ as a function of similarity $s(x)$ using the logistic family. We require the following structural properties from $\mathcal{L}(s(x),t,\gamma)$: 1) monotonicity in similarity, 2) boundedness in $[0,1]$, and 3) a low-dimensional parameterization so that we can obtain tight confidence bands for $(t,\gamma)$ from the relatively small observation sets $\mathcal{O}_{nn(x)}$. The logistic sigmoid is a canonical choice satisfying these properties and allows efficient MLE.
To empirically validate this choice, we analyze the SemBenchmarkLmArena configuration with text-embedding-3-small and GPT-4.1-nano. We randomly select and evaluation run and for each cached embedding we collect all observed pairs $(s(x),c(x))\in\mathcal{O}_{nn(x)}$, where $s(x)$ is the similarity between $x$ and $nn(x)$ and $c(x)\in\{0,1\}$ indicates whether returning the cached response was correct. Figure 7 shows, for the two embeddings with the largest numbers of labeled observations (42 and 33, respectively), a 1D $k$-NN estimate ($k=5$) of $\Pr(c=1\mid s,nn(x))$ as a function of similarity $s$. Both curves are monotone and exhibit a S-shaped transition from low to high correctness as similarity increases, which is the behavior that Eq. 9 is designed to capture for each $nn(x)$ and then feed into $\tau_{nn(x)}(s(x))$ and $\hat{\tau}$ in Eq. 9 and Eq. 11.
Figure 7: Empirical probability of correctness as a function of similarity $s$ for the two cached embeddings with the largest numbers of labeled observations (42 and 33) in SemBenchmarkLmArena with text-embedding-3-small and GPT-4.1-nano. We estimate $\Pr(c=1\mid s,nn(x))$ using a 1D $k$-NN smoother with $k=5$. In both cases the curve is monotone and approximately sigmoidal in $s$, supporting the per-embedding sigmoid model $\mathcal{L}(s(x),t,\gamma)$ in Eq. 9.
Figure 8 shows an aggregate view across all cached embeddings in the same configuration. We bin all $(s(x),c(x))$ pairs by similarity and plot the empirical correctness probability $\Pr(c=1\mid s)$ per bin. The resulting curve is again monotone and clearly S-shaped, with low correctness at small similarity, a rapid increase in a mid-range similarity region, and saturation near one at high similarity.
Figure 8: Aggregate empirical correctness probability as a function of similarity $s$ over all cached embeddings in SemBenchmarkLmArena with text-embedding-3-small and GPT-4.1-nano. We bin all $(s(x),c(x))$ pairs by similarity and plot the fraction of correct reuse decisions per bin. The resulting curve is monotone and clearly S-shaped, with low correctness at low similarity, a sharp increase in a mid-similarity region, and saturation at high similarity, which matches the qualitative behavior modeled by the sigmoid family in Eq. 9.
Appendix F Additional Accuracy, Latency, and Baseline Evaluations
F.1 SemCacheSearchQueries Benchmark
@@
We evaluate whether vCache satisfies the user-defined error rate $\delta$ while maintaining competitive performance. As shown in Figure 4, vCache consistently remains below the specified error bound across all tested $\delta$ values. Moreover, as the error rate stabilizes, vCache continues to improve its cache hit rate (Figure 4, right). In contrast, GPTCache exhibits increasing error rates as the sample size grows, despite improved hit rates. This trend reflects a fundamental limitation of static thresholds: maintaining a bounded error rate requires continuously increasing the threshold. Over time, no static threshold below 1.0 may suffice to satisfy a strict error constraint, making such systems difficult to tune and unreliable at scale.
@@
F.2 SemCacheCombo Benchmark
vCache respects user-defined error-rate requirements on SemCacheCombo
Figure 4 evaluates whether vCache satisfies the user-defined error rate $\delta$ on the SemCacheCombo benchmark. Across all tested $\delta$ values, the realized error of vCache remains below the requested bound, confirming that the learned thresholds reliably enforce the target error rate. As more samples are processed, vCache maintains its empirical error while steadily increasing cache hit rate, indicating ongoing learning from additional data. In contrast, GPTCache with a fixed similarity threshold of $0.83$ exhibits growth in both error rate and hit rate as the sample size increases.
Dynamic, embedding-specific thresholds yield better trade-offs than static thresholds on SemCacheCombo
Figure 5 summarizes the resulting trade-offs by varying vCache’s error bound $\delta$ and GPTCache’s static similarity threshold. Each point corresponds to a full evaluation on SemCacheCombo, enabling a direct Pareto comparison. vCache traces a strictly better frontier: it achieves higher cache hit rates at the same error level, and lower error for comparable hit rates and latency. For example, vCache attains up to $12.5\times$ higher cache hit rates than the best static-threshold configuration while still satisfying the user-defined error bound. At matched average latency, vCache consistently delivers lower error than GPTCache. These results show that dynamic, embedding-specific thresholds provide superior accuracy–efficiency trade-offs to static thresholds.
F.3 SemCacheLMArena with OpenAI Embeddings
vCache respects user-defined error-rate requirements on SemCacheLMArena
Figure 4 evaluates whether vCache satisfies the user-defined error rate $\delta$ on the SemCacheLMArena benchmark. Across all tested $\delta$ values, the realized error rate of vCache remains below the user-defined target, confirming that the learned thresholds enforce the desired bound. As the sample size grows, vCache further reduces its empirical error while steadily increasing cache hit rate, indicating that it continues to learn from additional data. In contrast, GPTCache with a fixed similarity threshold of $0.98$ exhibits growth in both error rate and hit rate as more prompts arrive. Figure 5 outlines a Pareto comparison across all feasible thresholds.
Dynamic, embedding-specific thresholds yield better trade-offs than static thresholds
Figure 5 summarizes the overall accuracy–efficiency trade-offs on SemCacheLMArena by varying vCache’s error bound $\delta$ and GPTCache’s static similarity threshold. Each point represents a complete evaluation run, enabling a direct Pareto comparison. vCache traces out a strictly better frontier: it achieves substantially higher cache hit rates at the same error level and much lower error for comparable hit rates and latency. For example, vCache reaches a cache hit rate of $57\%$ while keeping the error rate below $0.5\%$, whereas GPTCache requires an order-of-magnitude higher error to obtain similar hit rates. At matched latency, vCache achieves up to $26\times$ lower error than the best static configuration. These outcomes demonstrate that dynamic, embedding-specific thresholds are better suited than static thresholds for maintaining strict error guarantees while exploiting semantic redundancy.
F.4 $\tau$ Computation Overhead
Figure 9 reports the latency of computing $\tau$ (Algorithm 1) on the SemCacheLMArena benchmark as a function of the current sample size. Each point corresponds to one update step, and the solid red line is a linear regression fit. The points form a tight horizontal band and the regression slope is effectively zero ($\approx 5.5\times 10^{-10}$ sec per sample), indicating that $\tau$ can be computed in constant time. Across all sample sizes, the latency remains below 1.5~1.5 ms, indicating that the overhead of computing $\tau$ is negligible.
Figure 9: Empirical latency of computing $\tau$ in Algorithm 1 as a function of sample size. The fitted regression line (slope $\approx 5.5\times 10^{-10}$) shows that the cost is effectively constant, adding at most $\approx 1.5$ ms (<0.0015<0.0015 s) per update.
F.5 Embedding Generation Overhead
@@
Figure 10: Embedding computation latency distributions across models, shown as 95th percentile whisker plots. GTE_FT = GteLargeENv1-5 (Zhang et al., 2024) fine-tuned (Zhu et al., 2024). E5 Large v2 FT = E5-large-v2 (Wang et al., 2022) fine-tuned (Zhu et al., 2024).
@@
F.6 Logistic Regression Latency Overhead
@@
F.7 vCache in combination with vLLM
@@
Baseline Config Overall Duration Avg. LLM Inference Latency Avg. Emb. Latency Cache Hit Rate Error Rate
@@
vLLM + vCache $\delta=0.01$ 214 min 0.32 sec 0.018 sec 18.1% 0.4%
@@
F.8 Additional Baseline Evaluations
@@
Baseline Threshold/ Delta Error Rate Cache hit Rate Comments and Observations
@@ F.9 Threshold Dilemma
Appendix G Hyperparameter Guidance
Appendix H Benchmark Creation
To the best of our knowledge, no open-source benchmarks currently exist for evaluating the performance and applicability of semantic caching systems. To address this gap, we construct and release four diverse benchmarks [^5] https://huggingface.co/vCache, each designed to reflect a distinct real-world use case: classification tasks, conversational chatbots, and search engines. This section describes the motivation and construction process behind each benchmark.
H.1 SemCacheClassification Benchmark
@@ H.3 SemCacheSearchQueries Benchmark
H.4 SemCacheCombo Benchmark
We introduce the SemCacheCombo benchmark to evaluate semantic caching in workloads that contain both reusable responses and unique, non-reusable responses. The dataset is constructed by combining two sources into a single sequence of 27,500 prompts. First, we take the SemCacheLMArena benchmark and select one representative prompt from each of its 3,500 semantic classes. Because each prompt represents a different cluster, these 3,500 queries are pairwise semantically distinct and should not share a reusable response. From the perspective of a semantic cache, every SemCacheLMArena prompt is therefore expected to result in a cache miss; any cache hit on this subset is, by definition, an incorrect reuse. Second, we append 24,000 prompts from the SemCacheClassification benchmark. Correct cache hits may only occur within this set of prompts. Next, we randomly shuffle all 27,500 prompts.